# Bayesian Optimisation with Unknown Hyperparameters: Regret Bounds Logarithmically Closer to Optimal

### v1

@@ Appendix D Unknown RKHS norm
$v(i)=B_{0}e^{i}.$
For RKHS being selected by the candidate-suggesting function of Definition D.1 and length scale being selected by the one of Definition 3.1, we get:
Lemma D.1.
In case of both RBF and $\nu$-Matérn kernel we have that:
@@
$\frac{R^{(\hat{\theta},B(\hat{\theta},\hat{N}))}(T)}{R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)}=\frac{C\sqrt{T\gamma_{T}^{\hat{\theta}}}\left(\sqrt{\gamma_{T}^{\hat{\theta}}}+B(\hat{\theta},\hat{N})\right)}{C\sqrt{T\gamma_{T}^{\theta^{\star}}}\left(\sqrt{\gamma_{T}^{\theta^{\star}}}+B^{\star}\right)}\leq 2\left(\frac{\theta^{\star}}{\hat{\theta}}\right)^{\hat{d}}\leq 2\left(\frac{q(i^{\star})}{q(i^{\star}+1)}\right)^{\hat{d}}=2\left(e^{-1/\hat{d}}\right)^{\hat{d}}=\mathcal{O}(1),$
where $i^{\star}=\max\{i\in\mathbb{N}\mid q(i)\geq\theta^{\star}\}$ and as such we have $\theta^{\star}\leq q(i^{\star})$ and $q(i^{\star}+1)\leq\hat{\theta}$.
@@
$\frac{R^{(\hat{\theta},B(\hat{\theta},\hat{N}))}(T)}{R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)}$ $=\frac{C\sqrt{T\gamma_{T}^{\hat{\theta}}}\left(\sqrt{\gamma_{T}^{\hat{\theta}}}+B(\hat{\theta},\hat{N})\right)}{C\sqrt{T\gamma_{T}^{\theta^{\star}}}\left(\sqrt{\gamma_{T}^{\theta^{\star}}}+B^{\star}\right)}\leq\frac{2\sqrt{\gamma_{T}^{\hat{\theta}}}B(\hat{\theta},\hat{N})}{\sqrt{\gamma_{T}^{\theta^{\star}}}B^{\star}}\leq 2\left(\frac{q(i^{\star})}{q(i^{\star}+1)}\right)^{\hat{d}}\frac{v(j^{\star}+1)}{v(j^{\star})}$
$\leq 2\left(e^{-1/\hat{d}}\right)^{\hat{d}}e=2=\mathcal{O}(1)$
@@
12:        $U_{t+1}=\left\{u\in U_{t}:L_{t}(u)+\frac{1}{|S_{t}^{u}|}\sum_{\tau\in S_{t}^{u}}\beta_{\tau}^{u}\sigma^{u}_{\tau-1}(\bm{x}_{\tau})\geq\max_{u^{\prime}\in U_{t}}L_{t}(u^{\prime})\right\}$
@@
Theorem D.1.
Let us use confidence parameters of $\xi_{t}=2\sigma_{N}^{2}\log(\hat{d}\ln g(T)\log b(T)\pi^{2}t^{2})-\log(3\delta)$ and $\beta^{\theta,B}_{t}=(\theta_{0}/\theta)^{d/2}B+\sigma_{N}\sqrt{2(\gamma^{\theta}_{t-1}+1+\ln(2/\delta))}$, then Algorithm 3 achieves with probability at least $1-\delta$ the cumulative regret $R_{T}$ of the algorithm admits the following bound:
$(\displaystyle R_{T}=\mathcal{O}\Bigg{(}$$\left(t_{0}+\iota\right)B(\theta^{\star},N^{\star})+$
$\left(R^{(\theta^{*},B(\theta^{\star},N^{\star}))}(T)+\sqrt{T\xi_{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{*}}\right)^{\hat{d}/2}\max\left(\left(\frac{\theta_{0}}{\theta^{*}}\right)^{\hat{d}/2},\frac{B(\theta^{\star},N^{\star})}{B_{0}}\right)\sqrt{\iota}+\iota\right)\Bigg{)},$
where $t_{0}=\max\{g^{-1}\left(e^{-1/\hat{d}}\theta_{0}/\theta^{\star}\right),b^{-1}(N^{\star}/B_{0})\}$ and $\iota=\hat{d}\ln g(T)\log b(T)$.
@@
Similarily as in the proof of Theorem 4.1, we look for $t_{0}$, such that at least one well-specified hyperparameter value will enter the considered set. This happens after at most $t_{0}=\max\{g^{-1}(\frac{\theta_{0}}{\theta^{\star}e^{1/\hat{d}}}),b^{-1}(\frac{B}{B_{0}}e)\}$. Observe that the set of all hyperparameters introduced by time step $T$ is $A=\Theta_{T}\times\mathcal{B}_{t}$. We thus have:
@@
We now apply Lemma A.1 to get:
$\sum_{t\notin\mathcal{T}}r_{t}\leq\mathcal{O}\left(|A|B+\left(R^{u^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\sqrt{\sum_{u\in A}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}+|A|\right)\right),$
where now $|A|=|\Theta_{T}||\mathcal{B}_{T}|=q^{-1}(\frac{\theta_{0}}{g(T)})v^{-1}(B_{0}g(T))=\hat{d}\log(g(T))\log(b(T))$. We now derive a similar Lemma as Lemma C.1.
Lemma D.2.
For the same type of kernels with different length scales, we have that
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq\sqrt{\left(\frac{\theta_{0}}{\theta^{*}}\right)^{\hat{d}}}\max\left(\sqrt{\left(\frac{\theta_{0}}{\theta^{*}}\right)^{\hat{d}}},\frac{N^{\star}}{B_{0}}\right)$
, where $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ in case of an $\nu$-Matern kernel.
@@
If $|S_{t}^{u^{*}}|\geq|S_{t}^{u}|$, the bound holds trivially. Thus we will assume $|S_{t}^{u^{*}}|<|S_{t}^{u}|$, which implies $\gamma_{|S_{t}^{u^{*}}|}\geq\gamma_{|S_{t}^{u}|}$ and $u\geq u^{\star}$. The suspected regret bounds are of the form:
$R^{u}(t)=C\sqrt{T\gamma_{T}^{u}}\left(\sqrt{\gamma_{T}^{u}}+B^{u}\right).$
@@
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq 2\frac{\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{u}}\left(\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{u}}+B^{u^{\star}}\right)}{\sqrt{\gamma_{|S_{t}^{u}|}^{u}}\left(\sqrt{\gamma_{|S_{t}^{u}|}^{u}}+B^{u}\right)}$
Case 1: Consider the case when $\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{u^{\star}}}\geq B^{u^{\star}}$. We then have:
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq 2\frac{\gamma_{|S_{t}^{u^{*}}|}^{u}}{\gamma_{|S_{t}^{u}|}^{u}}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}}\frac{\gamma_{|S_{t}^{u}|}^{u}}{\gamma_{|S_{t}^{u}|}^{u}}=2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}}.$
Case 2 Consider the case when $\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{u}}<B^{u}$. We then have:
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq 2\frac{\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{u}}B^{u^{\star}}}{\sqrt{\gamma_{|S_{t}^{u}|}^{u}}B^{u}}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}/2}\sqrt{\frac{\gamma_{|S_{t}^{u}|}^{u}}{\gamma_{|S_{t}^{u}|}^{u}}}\frac{B^{u^{\star}}}{B^{u}}=2\left(\frac{\theta}{\theta^{\star}}\right)^{\hat{d}/2}\frac{B^{u^{\star}}}{B^{u}}.$
Plugging expression for $|A|$ and using Lemmas D.1 and D.2 finishes the proof.
Appendix E Experiments Details

### v3

@@ Appendix F Unknown RKHS norm
$v(i)=N_{0}e^{i}.$
For RKHS being selected by the candidate-suggesting function of Definition F.1 and length scale being selected by the one of Definition 3.1, we get:
Lemma F.1.
In the case of both RBF and $\nu$-Matérn kernel, we have that:
@@
$\frac{R^{(\hat{\theta},B(\hat{\theta},\hat{N}))}(T)}{R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)}=\frac{\sqrt{T\gamma_{T}^{\hat{\theta}}}\left(\sqrt{\gamma_{T}^{\hat{\theta}}}+B(\hat{\theta},\hat{N})\right)}{\sqrt{T\gamma_{T}^{\theta^{\star}}}\left(\sqrt{\gamma_{T}^{\theta^{\star}}}+B(\theta^{\star},N^{\star})\right)}\leq 2\left(\frac{\theta^{\star}}{\hat{\theta}}\right)^{d}\leq 2\left(\frac{q(i^{\star})}{q(i^{\star}+1)}\right)^{d}=2\left(e^{-1/d}\right)^{d}=\mathcal{O}(1),$
where $i^{\star}=\max\{i\in\mathbb{N}\mid q(i)\geq\theta^{\star}\}$ and as such we have $q(i^{\star}+1)=\hat{\theta}\leq\theta^{\star}\leq q(i^{\star})r$.
@@
$\frac{R^{(\hat{\theta},B(\hat{\theta},\hat{N}))}(T)}{R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)}$ $=\frac{\sqrt{T\gamma_{T}^{\hat{\theta}}}\left(\sqrt{\gamma_{T}^{\hat{\theta}}}+B(\hat{\theta},\hat{N})\right)}{\sqrt{T\gamma_{T}^{\theta^{\star}}}\left(\sqrt{\gamma_{T}^{\theta^{\star}}}+B(\theta^{\star},N^{\star})\right)}\leq\frac{2\sqrt{\gamma_{T}^{\hat{\theta}}}B(\hat{\theta},\hat{N})}{\sqrt{\gamma_{T}^{\theta^{\star}}}B(\theta^{\star},N^{\star})}\leq 2\left(\frac{q(i^{\star})}{q(i^{\star}+1)}\right)^{d}\frac{v(j^{\star}+1)}{v(j^{\star})}$
$\leq 2\left(e^{-1/d}\right)^{d}e=2=\mathcal{O}(1)$
where $i^{\star}=\max\{i\in\mathbb{N}\mid q(i)\geq\theta^{\star}\}$ and $j^{\star}=\max\{j\in\mathbb{N}\mid v(j)\leq N^{\star}\}$ as such we have $q(i^{\star}+1)=\hat{\theta}\leq\theta^{\star}\leq q(i^{\star})$ and $v(j^{\star}+1)=\hat{N}\geq N^{\star}\geq v(j^{\star})$. ∎
@@
12:        $U_{t+1}=\left\{u\in U_{t}:L_{t}(u)+\frac{2}{|S_{t}^{u}|}\sum_{\tau\in S_{t}^{u}}\beta_{\tau}^{u}\sigma^{u}_{\tau-1}(\bm{x}_{\tau})\geq\max_{u^{\prime}\in U_{t}}L_{t}(u^{\prime})\right\}$
@@
Theorem F.1.
Let us use confidence parameters of $\xi_{t}=2\sigma_{N}^{2}\log(d\ln g(T)\log b(T)\pi^{2}t^{2})-\log(3\delta)$ and $\beta^{\theta,N}_{t}=B(\theta,N)+\sigma_{N}\sqrt{2(\gamma^{\theta}_{t-1}+1+\ln(2/\delta))}$, then Algorithm 3 achieves with probability at least $1-\delta$ the cumulative regret $R_{T}$ of the algorithm admits the following bound:
$(\displaystyle R_{T}=\mathcal{O}\Bigg{(}$$\left(t_{0}+\iota\right)B^{\star}+$
$\left(R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)+\sqrt{T\xi_{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}\frac{N^{\star}}{N_{0}}d\ln\frac{\theta_{0}}{\theta^{\star}}\ln\frac{N^{\star}}{N_{0}}+\iota\right)\Bigg{)},$
where $t_{0}=\max\{g^{-1}\left(e^{-1/d}\theta_{0}/\theta^{\star}\right),b^{-1}(N^{\star}/N_{0})\}$ and $\iota=d\ln g(T)\log b(T)$.
@@
Similarly as in the proof of Theorem 4.1, we look for $t_{0}$, such that at least one well-specified hyperparameter value will enter the considered set. This happens after at most $t_{0}=\max\{g^{-1}(\frac{\theta_{0}}{\theta^{\star}e^{1/d}}),b^{-1}(\frac{B}{B_{0}}e)\}$. Observe that the set of all hyperparameters introduced by time step $T$ is $A=\Theta_{T}\times\mathcal{B}_{t}$. We thus have:
@@
We now apply Lemma B.5 to get:
$\sum_{t\notin\mathcal{T}}r_{t}\leq\mathcal{O}\left(|A|B+\left(R^{u^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\sum_{u\in\mathcal{M}_{0}}\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+|A|\right)\right),$
where now $|A|=|\Theta_{T}||\mathcal{B}_{T}|=q^{-1}(\frac{\theta_{0}}{g(T)})v^{-1}(N_{0}b(T))=d\log(g(T))\log(b(T))$. We now derive a Lemma similar to Lemma D.1.
Lemma F.2.
If the event of Lemma B.5 holds, then for any $\theta\in\mathcal{M}_{0}$ and $t\geq t_{0}$ we have that
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}\frac{N^{\star}}{N_{0}}.$
Plugging expression for $|A|$, using Lemmas F.1 and F.2 and the fact that $|\mathcal{M}_{0}|=\mathcal{O}(d\ln\theta^{\star}\ln N^{\star})$ finishes the proof.
Appendix G Proof of Lemma F.2
See F.2
@@
If $|S_{t}^{u^{*}}|\geq|S_{t}^{u}|$, the bound holds trivially. Thus we will assume $|S_{t}^{u^{*}}|<|S_{t}^{u}|$. The suspected regret bounds are of the form:
$R^{u}(t)=\sqrt{T\gamma_{T}^{\theta}}\left(\sqrt{\gamma_{T}^{\theta}}+B(\theta,N)\right).$
@@
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq 2\frac{\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{\theta^{\star}}}\left(\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{\theta^{\star}}}+B(\theta^{\star},N^{\star})\right)}{\sqrt{\gamma_{|S_{t}^{u}|}^{\theta}}\left(\sqrt{\gamma_{|S_{t}^{u}|}^{\theta}}+B(\theta,N)\right)}$
Case 1: Consider the case when $\sqrt{\gamma_{|S_{t}^{\theta^{\star}}|}^{\theta^{\star}}}\geq B(\theta^{\star},N^{\star})$. We then have:
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\leq 2\frac{\gamma_{|S_{t}^{u^{*}}|}^{\theta^{\star}}}{\gamma_{|S_{t}^{u}|}^{\theta}}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{d}\frac{\gamma_{|S_{t}^{u}|}^{\theta}}{\gamma_{|S_{t}^{u}|}^{\theta}}=2\left(\frac{\theta}{\theta^{\star}}\right)^{d}\leq 2\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}\frac{N^{\star}}{N_{0}}.$
Case 2 Consider the case when $\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{\theta^{\star}}}<B(\theta^{\star},N^{\star})$. We then have:
$\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}$ $\leq 2\frac{\sqrt{\gamma_{|S_{t}^{u^{*}}|}^{\theta^{\star}}}B(\theta^{\star},N^{\star})}{\sqrt{\gamma_{|S_{t}^{u}|}^{\theta}}B(\theta,N)}\leq 2\left(\frac{\theta}{\theta^{\star}}\right)^{d/2}\sqrt{\frac{\gamma_{|S_{t}^{u}|}^{\theta^{\star}}}{\gamma_{|S_{t}^{u}|}^{\theta}}}\frac{B(\theta^{\star},N^{\star})}{B(\theta,N)}=2\left(\frac{\theta}{\theta^{\star}}\right)^{d}\frac{N^{\star}}{N}$
$\leq 2\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}\frac{N^{\star}}{N_{0}}.$
Appendix H Derivation of optimality rates
To obtain rates for A-GP-UCB, we use Corrolary 3 of [6]. While A-GP-UCB considered the case of unknown norm and bound simultaneously, to obtain the rate for unknown length scale only, we ignore the growth function used for the norm. Note that since, for A-GP-UCB $R_{T}=\mathcal{O}(b(T)g(T)^{d}R^{u^{\star}}(T))$ and in BO $R^{u}(T)=\sqrt{T\gamma_{T}^{u}}(\sqrt{B^{u}}+\sqrt{\gamma^{u}_{T}})$, if $b(T)g(T)^{d}$ grows at least as fast as $\sqrt{TB}$, then bound on $R_{T}$ grows at least as fast as $B^{u^{\star}}T$ and becomes trivial. Thus for the regret bound of A-GP-UCB to be meaningful, we have to assume $b(T)g(T)^{d}$ grows slower than $\sqrt{TB}$.
Inspecting the bounds of LB-GP-UCB and LNB-GP-UCB in Theorems 4.1 and F.1, we see that the term with $R^{\theta^{\star}}(T)$ or $R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)$ will dominate the bound. This is because by the previous assumption on the growth of $b(T)g(T)^{d}$, we get that $\iota=\mathcal{O}(\ln b(T)d\ln g(T))\leq\mathcal{O}(\ln(BT))$ and $\sqrt{T\xi_{t}}=\mathcal{O}(\sqrt{T\log\ln b(T)d\ln g(T)})=\mathcal{O}(\sqrt{T\log\log TB})$ and in both RBF and $\nu$-Matérn cases regret bound grows at least as fast as $\sqrt{T\log T}$. Also the term $\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}\frac{N^{\star}}{N_{0}}d\ln\frac{\theta_{0}}{\theta^{\star}}\ln\frac{N^{\star}}{N_{0}}$ is a constant and will eventually get dominated by $\iota$. We thus get that the bound will become dominated by $\iota R^{\theta^{\star}}(T)$ or $\iota R^{(\theta^{\star},B(\theta^{\star},N^{\star}))}(T)$ and the suboptimality is just $\iota$.
Appendix I Experiments Details